OpenStudy (anonymous):

Determine whether each equation defines y as a function of x 1) |y|=4-x 2) -x+y^2=5 3) xy=8 4) x+y^3=9

OpenStudy (anonymous):

Isolate for y and see if you get unique answers.

OpenStudy (anonymous):

So basically see if x has more than one value for y, if it does, than it's not a function of y.

OpenStudy (anonymous):

so the first on we will have two answers because of the absolute value right

OpenStudy (anonymous):

one*

OpenStudy (anonymous):

y = 4-x y = x - 4

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

alright the second one would be like that right y^2=x+5 right

OpenStudy (anonymous):

is that gonna be the only answer or should i simplify more

OpenStudy (anonymous):

That will give you two solutions as well

OpenStudy (anonymous):

is it because of the power

OpenStudy (anonymous):

You have to simplify till you can isolate y on its own.

OpenStudy (anonymous):

so the third one is one solution because it would be like that y=8-x

OpenStudy (anonymous):

It's 8/x but yes, one solution :)

OpenStudy (anonymous):

oh right y=8/x

OpenStudy (anonymous):

and how a bout that one x+y^3=9

OpenStudy (anonymous):

Cube root :d

OpenStudy (anonymous):

so its going to be more than one solution

OpenStudy (anonymous):

I'd say so, but I hope you understand the question and what you have to do :)

OpenStudy (anonymous):

sure yah! but am just making sure about what i have just answered

OpenStudy (anonymous):

thank you so much for helping

OpenStudy (anonymous):

Np, take care :)